Engineering Design Process

Yousef Haik and Namas Chandra

Florida A& M University - Florida State University

College of Engineering Department of Mechanical Engineering

Tallahassee, FL 32310

Preface

Early human history shows that artifact design was based on human need. A knife was invented to protect human from his surrounding. Egyptian invented the wheels to make their traveling and transportation easier and so on. The terminology ”the need is mother invention” was demonstrated by the different human inventions. The idea of design and development is what distinguishes engineering from science. Engineers throughout history have wrestled with problems of water not being where it was needed, of minerals not being close at hand, of building materials having to be moved. Ancient engineers were often called upon to devise means of erecting great monuments, to design defenses against enemies and to move people and goods across rough terrain and rougher water. These artifacts were primarily inventive in nature and predominantly characterized by cut and try approaches. Until Galileo and Newton who laid the foundations for the analytical methods that students now learn in engineering schools. The Industrial Revolution brought a profileration of new machines and manufacturing techniques and this in turn provided an impetus for the growth of science and commerce on the international scale. The history of engineering is rich with stories of success triumphaling over failure.

In early 1900s it was common in American industry for inventions to be made by a Master Mechanic and then a draftsman was called to copy on paper what had been synthesized experimentally in the shop. The value of synthesizing on paper was soon realized since it was less costly and more efficient to erase rather than to remake parts in the shop. Recent trends have been to apply theory where appropriate in the process of mechanical design. However, in the total picture it is emphasized that all of the useful ingredient, such as various aspects of art, science, engineering, practical experience and ingenuity must be properly blended in the design process. A successful design is acheived when a logical procedure is followed to meet a specific need.

The procedure, called the design process, is similar to the scientific method with respect to a step by step routine. The design process encompasses the following activities, all of which must be completed

1. Idenftification of a need

2. Preoblem definition (Goals)

3. Market analysis (Search)

4. Specifications and constarins

5. Function analysis

6. Generating ideas

7. Evaluating alternatives

8. Analysis

9. Experiment

10. Marketing

In most times, designs are not accomplished by an engineer simply completing the design steps shown in the given order. The design process hold within its structure an iterative procedure, as the engineer proceeds through the steps, new information may be discovered and new objectives may be specified then the steps need to be revisited again. The more time and effort an engineer spends on the problem definition and understanding the need statement, the less frequent the number of iteration that may be needed.

In these notes we tried to provide the student with a reference for the course and present the material in a systematic easy to understand manner. Students are highly encouraged to submitt their comments and suggestions to the authors.

Contents

Preface i

1 Introduction 1 1.1 Definition of Engineering Design . . . . . . . . . . . . . . . . . 3 1.2 Design Levels . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.3 Design Process . . . . . . . . . . . . . . . . . . . . . . . . . . 4

1.3.1 Identification the customer need . . . . . . . . . . . . . 5 1.3.2 Definition of Goals . . . . . . . . . . . . . . . . . . . . 10 1.3.3 Market Analysis . . . . . . . . . . . . . . . . . . . . . . 17 1.3.4 Establishing Functions . . . . . . . . . . . . . . . . . . 18 1.3.5 Task Specifications . . . . . . . . . . . . . . . . . . . . 18 1.3.6 Conceptualization . . . . . . . . . . . . . . . . . . . . . 19 1.3.7 Evaluating Alternatives . . . . . . . . . . . . . . . . . 21 1.3.8 Analysis and Optimization . . . . . . . . . . . . . . . . 21 1.3.9 Experiment . . . . . . . . . . . . . . . . . . . . . . . . 21 1.3.10 Marketing . . . . . . . . . . . . . . . . . . . . . . . . . 22

1.4 Professionalism and Ethics . . . . . . . . . . . . . . . . . . . . 23 1.5 Problems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29

1.5.1 Team Activities . . . . . . . . . . . . . . . . . . . . . . 29 1.5.2 Individual Activities . . . . . . . . . . . . . . . . . . . 30

2 Prerequisites 32

2.1 Scheduling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 2.1.1 Gantt Chart . . . . . . . . . . . . . . . . . . . . . . . . 32 2.1.2 CPM/PERT . . . . . . . . . . . . . . . . . . . . . . . . 33 2.1.3 Definitions . . . . . . . . . . . . . . . . . . . . . . . . . 34 2.1.4 Network Development . . . . . . . . . . . . . . . . . . 35

2.2 Material Selection . . . . . . . . . . . . . . . . . . . . . . . . . 37 2.2.1 Material Classifications and Properties . . . . . . . . . 39 2.2.2 Material Selection Process . . . . . . . . . . . . . . . . 40 2.2.3 Primary Manufacturing Methods . . . . . . . . . . . . 41

2.3 Human Factors . . . . . . . . . . . . . . . . . . . . . . . . . . 42 2.3.1 Human Sensory Capabilities . . . . . . . . . . . . . . . 43 2.3.2 Anthropometric Data . . . . . . . . . . . . . . . . . . . 45

2.4 Safety Considerations . . . . . . . . . . . . . . . . . . . . . . . 45

2.4.1 Safety Analysis Techniques . . . . . . . . . . . . . . . . 45 2.5 Terchnical Writing and Presentation . . . . . . . . . . . . . . 47

2.5.1 Steps in Writing a Report . . . . . . . . . . . . . . . . 49 2.5.2 Illustration Guidelines . . . . . . . . . . . . . . . . . . 50 2.5.3 Mechanics of writing . . . . . . . . . . . . . . . . . . . 51

2.6 Presentation Style . . . . . . . . . . . . . . . . . . . . . . . . . 52 2.6.1 Objective . . . . . . . . . . . . . . . . . . . . . . . . . 52 2.6.2 Oral Presentation - Obstacles . . . . . . . . . . . . . . 52 2.6.3 Oral Presentation Dos and Donts . . . . . . . . . . . . 53 2.6.4 Oral Presentation Techniques . . . . . . . . . . . . . . 54 2.6.5 Question/Answer session . . . . . . . . . . . . . . . . . 55

2.7 Problems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 2.7.1 Team Activities . . . . . . . . . . . . . . . . . . . . . . 55 2.7.2 Indvidual Activities . . . . . . . . . . . . . . . . . . . . 57

3 Needs and Goals 63 3.1 Problem Definition: Need Statement . . . . . . . . . . . . . . 63 3.2 Objective Tree . . . . . . . . . . . . . . . . . . . . . . . . . . . 64

3.2.1 Worked Example . . . . . . . . . . . . . . . . . . . . . 65 3.2.2 Example: Automatic coffee maker . . . . . . . . . . . 66 3.2.3 Example: Automatic aluminum can crusher . . . . . . 68

3.3 Problems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 3.3.1 Team Activities . . . . . . . . . . . . . . . . . . . . . . 71

4 Market Analysis and Information Gathering 76 4.1 How to Conduct a Market Analysis . . . . . . . . . . . . . . . 77 4.2 Relevant Information Resources . . . . . . . . . . . . . . . . . 79

4.2.1 Industry Information . . . . . . . . . . . . . . . . . . 79 4.2.2 Market Information . . . . . . . . . . . . . . . . . . . 84

4.3 Web Tools . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 4.4 Example . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 4.4.1 Problem Definition . . . . . . . . . . . . . . . . . . . . 86 4.4.2 Potential cutomers . . . . . . . . . . . . . . . . . . . . 86 4.4.3 Companies . . . . . . . . . . . . . . . . . . . . . . . . . 87 4.4.4 SIC Code . . . . . . . . . . . . . . . . . . . . . . . . . 87 4.4.5 Trends . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 4.4.6 Market Infromation . . . . . . . . . . . . . . . . . . . . 88 4.4.7 Patents . . . . . . . . . . . . . . . . . . . . . . . . . . 88

4.5 Problems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 4.5.1 Team Activities . . . . . . . . . . . . . . . . . . . . . . 88 4.5.2 Individual Activities . . . . . . . . . . . . . . . . . . . 88

5 Establishing Functional Structure 89

5.1 Procedure . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94

5.1.1 Example of a washing machine . . . . . . . . . . . . . 96 5.2 Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98

5.2.1 Some examples of overall function . . . . . . . . . . . . 99 5.3 sub-function . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 5.4 Example . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 5.5 Problems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100

5.5.1 Team Activities . . . . . . . . . . . . . . . . . . . . . . 100 5.5.2 Individual Activities . . . . . . . . . . . . . . . . . . . 100

6 Specifications 102 6.1 Procedure for Performance Specification Method . . . . . . . . 105 6.2 Example: Automatic can crusher . . . . . . . . . . . . . . . . 110 6.3 Quality Function Deployment Method (QFD) . . . . . . . . . 110 6.4 Example:Automatic can crusher . . . . . . . . . . . . . . . . . 113 6.5 Problems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114

6.5.1 Team activities . . . . . . . . . . . . . . . . . . . . . . 114 6.5.2 Individual Activities . . . . . . . . . . . . . . . . . . . 114

7 Developing Concepts 123 7.1 Visual Thinking and Creativity . . . . . . . . . . . . . . . . . 124 7.2 Developing Working Structures . . . . . . . . . . . . . . . . . 126 7.3 Steps to Develop Concepts from Functions . . . . . . . . . . . 128 7.4 Sources of Concept Ideas: Techniques to Enhance Creativity . 130 7.5 EXAMPLES . . . . . . . . . . . . . . . . . . . . . . . . . . . . 132

7.5.1 Automatic can crusher . . . . . . . . . . . . . . . . . . 133 7.6 Problems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135

7.6.1 Team Activity . . . . . . . . . . . . . . . . . . . . . . . 135

8 Concepts Evaluation 137 8.1 Decision Matrix . . . . . . . . . . . . . . . . . . . . . . . . . . 138 8.2 Example . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 140 8.3 Example:Automatic can crusher . . . . . . . . . . . . . . . . . 142

9 Embodiment Design 148 9.1 Product Drawings . . . . . . . . . . . . . . . . . . . . . . . . . 149 9.2 Bill of Material . . . . . . . . . . . . . . . . . . . . . . . . . . 151 9.3 Prototype . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151 9.4 Example: Automatic can crusher . . . . . . . . . . . . . . . . 152 9.5 Problems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153

9.5.1 Team Activity . . . . . . . . . . . . . . . . . . . . . . . 153

10 Cost Analysis 157 10.1 Costs Classifications . . . . . . . . . . . . . . . . . . . . . . . 159 10.2 Cost Estimate Methods . . . . . . . . . . . . . . . . . . . . . . 161 10.3 Labor Costs . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 10.4 Product Pricing . . . . . . . . . . . . . . . . . . . . . . . . . . 163

10.4.1 Break-Even Chart . . . . . . . . . . . . . . . . . . . . . 163 10.4.2 Linear Programming . . . . . . . . . . . . . . . . . . . 165

10.5 Problems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168 10.5.1 Team Activities . . . . . . . . . . . . . . . . . . . . . . 168 10.5.2 individual activities . . . . . . . . . . . . . . . . . . . . 170

Chapter 1

Introduction

An engineering student during his training is crammed with a vast amount of theoretical material and information. He only realizes his helplessness when he is faced with the task of logically applying what he has learned to a specific end. So long as his work is based on familiar models or previous designs, the knowledge he possesses is perfectly adequate to enable him to find a solution on conventional lines. As soon as he is required to develop something already in existence to a more advanced stage, or to create something entirely new without a previous design to guide him, he will fail miserably unless he has reached a higher level of understanding. The design process was formalized to enable student in following a systematic approach in design. The most important step of the design process is identifying the customer need. Most of the time the customer provide a generic statement of need and it is to the engineer to identify the real need of the customer. lets consider an example to illustrate the point:

We are required to design a chair that can be used by a child?. Clearly all of us know how to sit on a chair, so in that prospectus we know how it works. A chair is used for sitting. unfortunately this description does not say how the chair is made?. What material is used?. Is the chair flexible or rigid?. Does the chair rotate or fixed?. What does it mean, it is to be used by a child?. Is the safety the biggest concern?. How much does it cost? How old is the child? etc.

Many factors in the engineering design does not have a mathematical model to be based on, but the process is maintained to be logically driven in systematic sequence. In the previous example we can use Newton’s Laws of equilibrium to describe the forces generated in the chair’s legs. We can also describe the deformation of the legs when a person sits down. We can even use the finite element analysis to estimate the stresses in the legs, the seat and the back. We can also describe the make and the joints used in the chair. Some of these indeed can be represented using mathematical models but what mathematical model will describe the color of the chair, what mathematical model will measure the how safe is the chair to be used by a child. It is clearly appearant that the function of the chair cannot be completely be presented by a mathematical model. These functions are manipulated by reasoning. It is clear that different schemes may be needed to describe a certain design such as analytical based models that can identify the geometric presentation of the artifact, economic schemes that describe the cost of producing the artifact, verbal contents that describe the function of the artifacts and so on. So it is quit clear that in the designing process the mathematical modeling although important but not sufficient to design the artifact.

The design engineer must learn to think independently, to draw conclusions and to combine. There are many who believe they can acquire all this by attending lectures and reading textbooks. What they fail to realize, is that they are only accumulating one fresh item of knowledge after another. Understanding, coupled with powers of logical deduction and judgment is not a capability that can be conferred from outside; on the contrary, it is something purely personal and inward acquired only by diligent thinking and working with the knowledge already possessed. A basic pre condition for independent designing is a lively imagination to do original work.

Through the history, human have been successfully designing artifacts to satisfy the needs of the civilization. The history is full of great designs and invention. Recently, design is driven to meet an existing completion, reducing a hazard or an inconvenience or developing a new approach. Not all what the engineer build has been successful, occasionally, catastrophic failures occur. Few of the well publicized disasters associated with engineering systems were:

• Chernobyle nuclear power plant

• Challenger space shuttle

• Boing 737 lost its cabin roof during flight in 1988

• A skywalk of Kansas City Hyatt Regency Hotel collapsed just after the hotel was opened in 1981.

• DC-10 airplane lost an engine in flight and crashed in 1979 Walton [8] lists the reason for failures in most engineering designs as

1. Incorrect or overextended assumptions

2. Poor understanding of the problem to be solved

3. Incorrect design specifications

4. Faulty manufacturing and assembly

5. Error in design calculations

6. Incomplete experimentation and inadequate data collection

7. Errors in drawings

8. Faulty reasoning from good assumption

1.1 Definition of Engineering Design

Dym [3] defines design as the systematic, intelligent generation and evaluation of specifications for artifacts whose form and function achieve stated objectives and satisfy specified constrains. A formal definition of engineering design is found in the curriculum guidelines of the Accreditation Board for Engineering and Technology (ABET). The ABET definitions states: Engineering design is the process of devising a system, component, or process to meet desired needs. It is a decision-making process (often iterative) , in which the basic sciences, mathematics, and engineering sciences are applied to convert resources optimally to meet a stated objective. Among the fundamental elements of the design process are the establishment of objectives and criteria, synthesis, analysis, construction, testing, and evaluation. The engineering design component of a curriculum must include most of the following features: development of student creativity, use of open-ended problems, development and use of modern design theory and methodology, formulation of design problem statement and specifications, production processes, concurrent engineering design, and detailed system description. Further, it is essential to include a variety of realistic constraints such as economic factors, sfatey, reliability, aesthetics, ethics and social impact.

1.2 Design Levels

As in any field of human activity there are different degrees of difficulty. In design they are adaptive designs, developed designs and new design.

1. Adaptive Design: In the great majority of instances the designer’s work will be concerned with adaptation of existing designs. There are branches of manufacturing where development has practically ceased, so that there is hardly anything left for the designer to do except make minor modifications, usually in the dimensions of the product. Design activity of this kind therefore demands no special knowledge or skill and the problems presented are easily solved by a designer with ordinary technical training.

2. Development Design: Considerably more scientific training and design ability are needed for development design. The designer starts from an existing design, but the final outcome may differ quite markedly from the initial product

3. New Design: Only a small number of designs are new designs.

1.3 Design Process

To design is to create new product that turn to profit and benefit society in some way. The ability to design requires both science and art. The science can be learned through a systematic process (outlined in this chapter), experience and problem solving technique (will be mastered during your college education). The art is gained by practice and a total involvement to become proficient. The design of a device or system can be done in one of two ways

1. Evolutionary change: When a product is allowed to evolve over a period of time with only slight improvement. This is done when there is no competition. The creative capabilities of the designer are limited

2. Innovation: The rapid scientific growth and technological discoveries plus competition among companies for their slice in the market has placed a great deal of emphasis on new products which draws heavily on innovations to produce marketable ideas. Here the creative skills and analytical ability of the design engineer plays an important role.